# Geodesic Motion of Test Particles around the Scalar Hairy Black Holes with Asymmetric Vacua

### v1

@@ I Introduction
So far GR has successfully passed numerous observational tests with remarkable precision since the groundbreaking proposal by A. Einstein at 1915, ranging from the perihelion shift of mercury at 1920, the discovery of the binary pulsars by R. Hulse and J. Taylor [31], and then the recent detection of gravitational waves from the merger of binary compact objects by LIGO-VIRGO-KAGRA (LVK) collaboration and imaging of shadow for supermassive black holes in the galaxies M87 and Sgr A by Event Horizon Telescope (EHT). Hence, these observational tests are actually the basic manifestation of GR about the notion of how the spacetime around a compact object can be bent by its own gravity and how does the corresponding spacetime affect the motion of a test particle, as summarized by the famous physicist J. Wheeler in his quotation “Matter tells space how to curve. Space tells matter how to move”. Therefore, we study the influence of nontrivial scalar field on the geodesic motion of test particles in the vicinity of HBH with asymmetric vacua [4, 5] and report the results in this paper. Besides, the geodesic motion of test particles in the vicinity of other types of black holes have been extensively studied, for instance, Schwarzschild black hole [32, 33], Schwarzschild-(A)dS black hole [34], swirling universe [35, 36, 37, 38], Myers-Perry black hole [39], regular black hole by Ayón-Beato and Gracía [40], black hole in braneworld [41], (2+1)-rotating black holes [42, 43], SU(2)-(A)dS black hole in conformal gravity [44], supersymmetric $\hbox{AdS}_{5}$ black hole [45], Kerr black hole [46, 47, 48, 49], Kerr-Newman black hole [50, 51, 52, 53, 54], $U(1)^{2}$ dyonic rotating black holes [55], Kerr-Newman-Taub-NUT black hole [56], black hole in the Einstein-Maxwell-dilaton-axion theory [57], polymer black hole in the loop quantum gravity [58], black hole in the Einsteinian cubic gravity [59], Carrollian Reissner-Nordstrom black hole [60]. The trajectories of test particles in these black holes can be described analytically by the elliptic functions, since the closed form of these black holes are known. However, we can only numerically obtain the trajectories of test particles in this paper since we couldn’t obtain the closed form of the HBH, which can only be constructed numerically by solving the Einstein equation.
@@ II Scalar Hairy Black Holes with Asymmetric Vacua
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{16\pi G}-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)\right]\,,$ (1)
@@
Figure 2: Two basic properties of HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$ when the scalar field $\phi_{H}$ is nontrivial at the horizon $r_{H}$ : (a) The reduced area of horizon $a_{H}$ vs $\phi_{H}$ (b) The reduced Hawking temperature vs $\phi_{H}$.
@@ III Geodesic Motion of Test Particles Around the HBH
The extremum of $v_{\text{eff}}(r)$ indicates the presence of a photon sphere, where the null geodesics can theoretically orbit arbitrary amount of time. Since the imaging of shadow for the supermassive black holes in the center of some galaxies can also be described by the framework of backward ray-tracing, therefore it might be convenient to show the trajectories of null geodesics around the HBH according to the number of intersections $n$ between the null geodesics and the accretion disk, which can be defined as
@@
This allows us to classify the trajectories of null geodesics into three types: direct emission with $40<n<3/$ 4 for $b\in\left(0,b_{2}^{-}\right)\cup\left(b_{2}^{+},\infty\right)$, lensing with 3/$43/4<n<5/$ 4 for $b\in\left(b_{2}^{-},b_{3}^{-}\right)\cup\left(b_{3}^{+},b_{2}^{+}\right)$ and photon sphere with $n>5/4$ for $b\in\left(b_{3}^{-},b_{3}^{+}\right)$. Note that $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$, where $m$ represents the number of intersections of null geodesics with the accretion disk and $b_{c}$ is the critical value of $b$ for the photon sphere.
@@ IV Results and Discussions
Figure 3: The effective potential $V_{\text{eff}}(x)$ in the compactified coordinate $x=1-r_{H}/r$ for a massive particle around the HBH with $r_{H}=1$ for a) $\phi_{1}=0.5$ and $\phi_{H}=0.3$; b) $\phi_{1}=0.5$ and $\phi_{H}=0.4999$; c) $\phi_{1}=1.0$ and $\phi_{H}=0.3$; d) $\phi_{1}=1.0$ and $\phi_{H}=0.995$; e) The bound orbit of a massive particle with $E^{2}=0.95$, labeled by “Orbit 1” in (a) around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$; f) A massive particle with $E^{2}=1.2$, labeled by “Orbit 2” in (a) fleeing away from the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$. Note that the black disc represents the HBH.
@@
Fig. 4 (a) and (b) shows the effective potential $v_{\text{eff}}(x)$ of the null geodesics in the compactified coordinate $x$ in the HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$, respectively. Both figures depict that $v_{\text{eff}}(x)$ possesses a local maximum, which indicates that the photon sphere is unstable. Recall that the location of photon sphere for the Schwarzchild black hole in the compactified coordinate $x$ is given by $x_{\text{ph}}=1/3$ [66], thus Fig. 4(c) shows the location of photon sphere for HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$ don’t deviate too much from the Schwarzschild black hole. Fig. 4 (d) compares the number of intersection of null geodesics with the accretion disk $n$ for the Schwarzschild black hole and the HBH with $\phi_{1}=0.5$. Although the trajectories of null geodesics around the HBH still can be divided into three types: direct, lensing and photon sphere, we observe that when $\phi_{H}$ increases, the values of impact parameters $b_{2}^{-}$, $b_{2}^{+}$, $b_{3}^{-}$ and $b_{3}^{+}$ associated to different types of trajectories as shown in Table 3 for the HBH can deviate significantly from the Schwarzcshild black hole. Besides, the three types of trajectories of null geodesics around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1.0$ can be visualized in Fig. 4 (e).
@@
Figure 4: The effective potential $v_{\text{eff}}(x)$ in the compactified coordinate $x=1-r_{H}/r$ for null geodesics around the HBH with $r_{H}=1$ for a) $\phi_{1}=0.5$; b) $\phi_{1}=1.0$; c) The location of photon sphere in the compactified coordinate $x$ around the HBH with $r_{H}=1$, $\phi_{1}=0.5$ and $\phi_{1}=1.0$ versus $\phi_{H}$; d) The number of intersection of null geodesics with accretion disk $n$ vs $\phi_{H}$ around the Schwarzschild black hole and HBHs; e) The trajectories of null geodesics consist of direct (black curves), lensing (yellow curves) and photon sphere (blue curve) around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$.
@@ V Conclusion and Outlook
For the photon around the HBHs with $r_{H}=1$ and $\phi_{1}=0.5,1.0$, its $V_{\text{eff}}(r)$ possesses a local maximum, which indicates the presence of unstable photon sphere. The location of photon sphere around the HBHs doesn’t deviate too much from the Schwarzschild black hole, perhaps the increment of $\phi_{1}$ could exhibit large deviation. According to the number of intersection of null geodesics with the accretion disk, the null geodesics can be divided into three types which are the direct, lensing and photon sphere with their associated impact parameters $b$. When the strength of scalar field at the horizon increases, the values of $b$ correspond to these three types of null geodesics can vary significantly from the Schwarzschild black hole.
Finally, several new research directions can be carried out from this work for the future research. First, we can continue to study the imaging of shadow with different emission profiles for this HBH, and compare them with the Schwarzschild black hole, since we have already calculated the null geodesics in this paper. Then, we can also study the imaging of shadow and then provide some constraints for some parameters for different HBHs [6, 8, 9] with different scalar potentials in the EKG theory.

### v2

@@ I Introduction
So far GR has successfully passed numerous observational tests with remarkable precision since the groundbreaking proposal by A. Einstein at 1915, ranging from the perihelion shift of mercury at 1920, the discovery of the binary pulsars by R. Hulse and J. Taylor [31], and then the recent detection of gravitational waves from the merger of binary compact objects by LIGO-VIRGO-KAGRA (LVK) collaboration and imaging of shadow for supermassive black holes in the galaxies M87 and Sgr A by Event Horizon Telescope (EHT). Hence, these observational tests are actually the basic manifestation of GR about the notion of how the spacetime around a compact object can be bent by its own gravity and how does the corresponding spacetime affect the motion of a test particle, as summarized by the famous physicist J. Wheeler in his quotation “Matter tells space how to curve. Space tells matter how to move”. Therefore, we study the influence of nontrivial scalar field on the geodesic motion of test particles in the vicinity of HBH with asymmetric vacua [4, 5] and report the results in this paper. Besides, the geodesic motion of test particles in the vicinity of other types of black holes have been extensively studied, for instance, Schwarzschild black hole [32, 33], Schwarzschild-(A)dS black hole [34], swirling universe [35, 36, 37, 38], Myers-Perry black hole [39], regular black hole by Ayón-Beato and Gracía [40], black hole in braneworld [41], (2+1)-rotating black holes [42, 43], SU(2)-(A)dS black hole in conformal gravity [44], supersymmetric $\hbox{AdS}_{5}$ black hole [45], Kerr black hole [46, 47, 48, 49], Kerr-Newman black hole [50, 51, 52, 53, 54], $U(1)^{2}$ dyonic rotating black holes [55], Kerr-Newman-Taub-NUT black hole [56], black hole in the Einstein-Maxwell-dilaton-axion theory [57], polymer black hole in the loop quantum gravity [58], black hole in the Einsteinian cubic gravity [59], Carrollian Reissner-Nordstrom black hole [60], Euclidean Schwarzschild geometry [61]. The trajectories of test particles in these black holes can be described analytically by the elliptic functions, since the closed form of these black holes are known. However, we can only numerically obtain the trajectories of test particles in this paper since we couldn’t obtain the closed form of the HBH, which can only be constructed numerically by solving the Einstein equation.
@@ II Scalar Hairy Black Holes with Asymmetric Vacua
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{16\pi G}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-V(\phi)\right]\,,$ (1)
@@
Figure 2: Two basic properties of HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$ when the scalar field $\phi_{H}$ is nontrivial at the horizon $r_{H}$ : (a) The reduced area of horizon $a_{H}$ vs $\phi_{H}$ (b) The reduced Hawking temperature vs $\phi_{H}$; (c) The parameter $\phi_{0}$ as the function of $\phi_{H}$ for $\phi_{1}=0.5$ and $\phi_{1}=1.0$; (d) The solutions $m(x)$, $\sigma(x)$ and $\phi(x)$ of the HBH with $\phi_{H}=0.7,\phi_{H}=1.0,r_{H}=1.0$ in the compactified coordinate $x$.
@@ III Geodesic Motion of Test Particles Around the HBH
The extremum of $v_{\text{eff}}(r)$ indicates the presence of a photon sphere, where the null geodesics can theoretically orbit arbitrary amount of time. Since the imaging of shadow for the supermassive black holes in the center of some galaxies can also be described by the framework of backward ray-tracing, therefore it might be convenient to show the trajectories of null geodesics around the HBH according to the total number of orbits of null geodesics $n$, which can be defined as [66]
@@
This allows us to classify the trajectories of null geodesics into three types: direct emission with $40<n<3/$ 4 for $b\in\left(0,b_{2}^{-}\right)\cup\left(b_{2}^{+},\infty\right)$ where the null geodesics only intersect with the accretion disk once $(m=1)$; lensing with 3/$43/4<n<5/$ 4 for $b\in\left(b_{2}^{-},b_{3}^{-}\right)\cup\left(b_{3}^{+},b_{2}^{+}\right)$ where the null geodesics intersect with the accretion disk twice $(m=2)$; and photon sphere with $n>5/4$ for $b\in\left(b_{3}^{-},b_{3}^{+}\right)$ where the null geodesics intersect with the accretion disk three times $(m=3)$. Note that $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$, where $m$ represents the number of intersections of null geodesics with the accretion disk and $b_{c}$ is the critical value of $b$ for the photon sphere.
@@ IV Results and Discussions
Figure 3: The effective potential $V_{\text{eff}}(x)$ in the compactified coordinate $x=1-r_{H}/r$ for a massive particle around the HBH with $r_{H}=1$ for a) $\phi_{1}=0.5$ and $\phi_{H}=0.3$. Note that in the figure, “Orbit 1” and “Orbit 2” refer to the motion of particle with $E^{2}=0.95$ in (e) and $E^{2}=1.2$ in (f), respectively; b) $\phi_{1}=0.5$ and $\phi_{H}=0.4999$; c) $\phi_{1}=1.0$ and $\phi_{H}=0.3$; d) $\phi_{1}=1.0$ and $\phi_{H}=0.995$; e) The bound orbit of a massive particle with $E^{2}=0.95$, labeled by “Orbit 1” in (a) around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$; f) A massive particle with $E^{2}=1.2$, labeled by “Orbit 2” in (a) fleeing away from the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$. Note that the black disc represents the HBH.
@@
Fig. 4 (a) and (b) shows the effective potential $v_{\text{eff}}(x)$ of the null geodesics in the compactified coordinate $x$ in the HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$, respectively. Both figures depict that $v_{\text{eff}}(x)$ possesses a local maximum, which indicates that the photon sphere is unstable. Recall that the location of photon sphere for the Schwarzchild black hole in the compactified coordinate $x$ is given by $x_{\text{ph}}=1/3$ [68], thus Fig. 4(c) shows the location of photon sphere for HBH with $\phi_{1}=0.5$ and $\phi_{1}=1.0$ don’t deviate too much from the Schwarzschild black hole. Fig. 4 (d) compares the total number of orbit $n$ for the Schwarzschild black hole and the HBH with $\phi_{1}=0.5$. Although the trajectories of null geodesics around the HBH still can be divided into three types: direct, lensing and photon sphere, we observe that when $\phi_{H}$ increases, the values of impact parameters $b_{2}^{-}$, $b_{2}^{+}$, $b_{3}^{-}$ and $b_{3}^{+}$ associated to different types of trajectories as shown in Table 3 for the HBH can increase significantly from the Schwarzcshild black hole, this can impact the brightness and size of the optical appearance (shadow) of the HBH. Besides, the three types of trajectories of null geodesics around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1.0$ can be visualized in Fig. 4 (e).
@@
Figure 4: The effective potential $v_{\text{eff}}(x)$ in the compactified coordinate $x=1-r_{H}/r$ for null geodesics around the HBH with $r_{H}=1$ for a) $\phi_{1}=0.5$; b) $\phi_{1}=1.0$; c) The location of photon sphere in the compactified coordinate $x$ around the HBH with $r_{H}=1$, $\phi_{1}=0.5$ and $\phi_{1}=1.0$ versus $\phi_{H}$; d) Based on $n=\varphi/(2\pi)$, the null geodesics around the Schwarzschild black hole with $r_{H}=1$ and HBH with $\phi_{1}=0.5$, $r_{H}=1$ can be categorized into three emissions as a function of impact parameter $b$: direct (black), lensing (yellow) and photon ring (red). e) The trajectories of null geodesics consist of direct (black curves), lensing (yellow curves) and photon sphere (blue curve) around the HBH with $\phi_{1}=0.5$, $\phi_{H}=0.3$, $r_{H}=1$.
@@ V Conclusion and Outlook
For the photon around the HBHs with $r_{H}=1$ and $\phi_{1}=0.5,1.0$, its $V_{\text{eff}}(r)$ possesses a local maximum, which indicates the presence of unstable photon sphere. The location of photon sphere around the HBHs doesn’t deviate too much from the Schwarzschild black hole, perhaps the increment of $\phi_{1}$ could exhibit large deviation. According to the total number of orbits, the null geodesics can be divided into three types which are the direct, lensing and photon sphere with their associated impact parameters $b$. When the strength of scalar field at the horizon increases, the values of $b$ correspond to these three types of null geodesics can vary significantly from the Schwarzschild black hole.
Finally, several new research directions can be carried out from this work for the future research. First, we can continue to study the imaging of shadow with different emission profiles for this HBH, and compare them with the Schwarzschild black hole, since we have already calculated the null geodesics in this paper. Recently, the shadow of HBH with inverted Higgs-like scalar potential [8] has been studied [69]. Then, we can also study the imaging of shadow and then provide some constraints for some parameters for different HBHs [6, 9] with different scalar potentials in the EKG theory.